Calculate the de Broglie wavelength of a proton moving at 3.10 ✕ 104 m/s and 2.07 ✕ 108 m/s.

(a)    3.10 ✕ 104 m/s
......... m

(b)    2.07 ✕ 108 m/s
........ m
Answer #1

Solution)

Part a)

De Broglie wavelength,

λ = h/(mv)

= 6.63*10^-34 / (1.672*10^-27 * 3.1*10^4)

=1.279*10^-11 m

===============

Part b)

Lambda =6.63*10^-34 / (1.672E-27 * 2.07E8)

=1.97*10^-15 m

=================